### v1

# Joint Pedestrian Trajectory Prediction through Posterior Sampling

Abstract
Joint pedestrian trajectory prediction has long grappled with the inherent unpredictability of human behaviors. Recent investigations employing variants of conditional diffusion models in trajectory prediction have exhibited notable success. Nevertheless, the heavy dependence on accurate historical data results in their vulnerability to noise disturbances and data incompleteness. To improve the robustness and reliability, we introduce the Guided Full Trajectory Diffuser (GFTD), a novel diffusion model framework that captures the joint full (historical and future) trajectory distribution. By learning from the full trajectory, GFTD can recover the noisy and missing data, hence improving the robustness. In addition, GFTD can adapt to data imperfections without additional training requirements, leveraging posterior sampling for reliable prediction and controllable generation. Our approach not only simplifies the prediction process but also enhances generalizability in scenarios with noise and incomplete inputs. Through rigorous experimental evaluation, GFTD exhibits superior performance in both trajectory prediction and controllable generation. See these results at https://sites.google.com/andrew.cmu.edu/posterior-sampling-prediction.
I INTRODUCTION
Pedestrian trajectory prediction is crucial for human-robot interaction systems such as autonomous driving, etc. The goal is to predict future paths based on previous movements and the environmental context of pedestrians. By accurately predicting pedestrian trajectories, autonomous systems can plan their actions accordingly, ensuring safe and efficient navigation in dynamic environments. However, the unpredictable and complicated nature of human behaviors makes this task challenging, especially with multiple agents involved where their interactions need to be considered.
To simplify the problem, early works [1, 2, 3, 4, 5] focused on marginal pedestrian trajectory prediction, which forecasts the path for each pedestrian independently. Such approaches require a downstream planning module to perform safety checks for every combination of the individual predictions. Even so, some combinations could still produce unrealistic self-collisions and lead to failure in challenging scenarios. As a result, joint pedestrian trajectory prediction, which predicts consistent trajectories for all agents together, has gained attention in the community. A joint metrics term is introduced by [6] into supervision loss, transforming the marginal predictor into a joint predictor. However, joint pedestrian trajectory prediction remains challenging since existing approaches heavily rely on accurate and complete historical data to incorporate temporal and social dynamics. This dependence results in their vulnerability to interference from noisy disturbances and missing data, significantly threatening their effectiveness in real-world applications, such as sensors under adverse weather conditions.
Figure 1: Illustration of existing trajectory prediction framework and our guided full trajectory diffuser framework. (a) Prior multi-agent trajectory prediction methods directly generate entire future trajectories through supervised learning. (b) Previous multi-agent trajectory prediction methods utilizing diffusion models generate future trajectories step-by-step during the denoising process. (c) Our method for multi-agent trajectory prediction incorporates history guidance within the diffusion framework. It predicts entire trajectories and supports additional capabilities in a single model, including Noise-Robust Trajectory Prediction, Incomplete Data Prediction, and Controllable Trajectory Generation.
To deal with the noises, previous efforts involved augmenting data with predefined noise and training models on noisy datasets [7], and it can be further improved by adversarial training procedures [8]. In addition, historical trajectories from sensors could be incomplete. Previous studies  [9, 10, 11, 12] proposed to reconstruct missing data and predict future trajectories during the training phase, requiring a specifically designed model and well-established training strategy. Such approaches optimize per-problem functions, lacking adaptability to diverse tasks across various contexts.
Draw inspiration from the diffusion model with its remarkable capability of capturing the complicated distribution, we propose a unified framework, named, Guided Full Trajectory Diffuser (GFTD) for the joint pedestrian trajectory prediction to better handle the disturbances and incompleteness. GFTD represents the entire trajectory distribution, both historical and future, with one diffusion model. We formulate trajectory prediction and controllable generation as inverse problems and solve them through posterior sampling techniques. Specifically, we sample in-distribution full trajectories based on historical trajectories and priors such as physical constraints and behavioral intentions. Without the necessity for explicit training to handle noisy and/or incomplete inputs, GFTD enables robust prediction and controllable generation—both achievable during the inference time. In a nutshell, our proposed framework streamlines the training process and offers adaptability to various scenarios at inference, providing a solution that can address all challenges without extra training requirements.
Our contributions can be summarized as follows:
(1) We introduce a novel permutation-invariant framework for representing the joint distribution of full trajectories (both historical and future), converting trajectory prediction and controllable generation into a unified inverse problem.
(2) We utilize posterior sampling to solve the formulated problem. With our approach, there is no need for specific treatments during the training phase, as it can generalize to various types of data imperfections solely at inference time with one trained model.
(3) Extensive experiments demonstrate that our model not only performs comparably in joint trajectory prediction but also excels in controllable generation, particularly in scenarios with noise injection and incomplete historical data.
II Related Work
II-A Pedestrian Trajectory Prediction
Pedestrian trajectory prediction is crucial for human-robot interaction systems, such as autonomous driving. This task involves forecasting the future movement paths of pedestrians, given their historical movements and the environment. However, it is challenging to predict their motion because of the diverse and unpredictable nature of human behaviors. To address this problem, two mainstream paradigms have been developed. The supervised learning approach [5] aims to minimize the differences between the ground truth trajectories and the predictions using L2 loss, etc. On the other hand, the generative learning approach [2] formulates the prediction as a task of generating conditional distributions of the future trajectory based on past trajectories. Among various generative models, the diffusion model [13] [14] has shown exceptional performance in pedestrian trajectory prediction. Through a conditional reverse diffusion process, diffusion model generates future trajectory distributions from a standard Gaussian distribution, which captures accurate and diverse predictions of future trajectories.
II-B Marginal and Joint Distribution
Recently, joint pedestrian trajectory prediction has gained significant attention. Traditional methods focus on marginal pedestrian trajectory prediction which treat each pedestrian independently, ignoring potential interactions and the impact on future paths. Instead, joint pedestrian trajectory prediction takes these interactions into consideration, aiming for more consistent and interconnected predictions. [6] incorporates joint metrics as the training objective, transforming the marginal trajectory predictor into a joint trajectory predictor. In this paper, we propose Guided Full Trajectory Diffuser (GTFD) for joint pedestrian trajectory prediction, leveraging the strengths of diffusion models in generating accurate and diverse future trajectory distributions while considering the interactions between pedestrians.
II-C Posterior Sampling for Inverse Problems
Posterior sampling is to infer the underlying distribution conditioned on the measurements, given the unconditional distribution. Previous works [15] [16] [17] have demonstrated great performance in general inverse problems of computer vision, such as image inpainting and denoising. Recent work involves physical constraints [18], the reward in reinforcement learning [19], behavior preference in traffic simulation [20] [21] and bias in trajectory prediction [22], extending a wide range of applications of posterior sampling. In this paper, we gain the insight that, in the task of trajectory prediction, historical trajectories can be regarded as observation or measurement. As a result, we can unleash the potential of posterior sampling in dealing with prediction under uncertainty.
III Methods
In this section, we undertake a comprehensive exposition of our novel multi-agent trajectory prediction framework, conceptualizing the prediction task through the lens of spatial-temporal inpainting. We will first introduce the problem formulation and preliminaries to diffusion models. Subsequently, we explain how such transformation is made possible and provide a detailed exposition of our novel prediction model within our framework.
Figure 2: Architecture of our proposed GFTD framework. During inference, GFTD samples from Gaussian noise, and iteratively recover data with the denoise module $\epsilon_{\theta}$. It takes in noisy latent node features $x_{t}^{i}$ and edge features $e$. After MLP encoding, the edge is augmented by concatenating the encoded node features and they are both sent into the Processor which consists of stacked Graph Attention (GAT) layers. We then map the nodes to the same dimensions as $x_{t}$ and add residual connection, resulting in the predicted intermediate noise $\epsilon_{t}$. After each denoise step, we address conditions through posterior sampling. Finally, the denoised latent nodes $x_{0}$ are converted to trajectory space.
III-A Preliminaries on Diffusion models
Diffusion models are a class of generative models that operate on the principle of stochastic processes. They define a forward diffusion process that corrupts data by progressively introducing Gaussian noise. Conversely, in the reverse process, diffusion models reconstruct data from noise by iteratively reducing the Gaussian noise.
According to Itô SDE, the forward data noising process is defined as the following form:
$dx=f(x,t)dt+g(t)dw,$ (1)
where $w$ is the standard Wiener process, $f$ is the drift coefficient, and $g$ is the diffusion coefficient.
The corresponding reverse SDE of Eq. (1) is defined as
$dx=[f(x,t)-g^{2}(t)\nabla_{x}{\log{p_{t}}(x)}]dt+g(t)dw,$ (2)
where $p_{t}(x)$ is the probability density at diffusion step $t$, and $\nabla_{x}{\log{p_{t}}(x)}$ is the time-dependent score function, which is approximated by training a neural network $s_{\theta}(x_{t},t)$ with score matching. Here, $x_{t}$ denotes the intermediate noisy data $x$ at denoise time $t$. The training objective is defined as
$\mathbb{E}_{t}\{\lambda(t)\mathbb{E}_{x_{0}}\mathbb{E}_{x_{t}|x_{0}}[\left\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}{\log{p}(x_{t}|x_{0})}\right\|_{2}^{2}]\},$ (3)
where $\lambda(t)$ is a weighting function, and $t$ is uniformly sampled over $[0,T]$.
In this paper, we follow the implementation introduced by Denoising Diffusion Probabilistic Models (DDPM) [23]. DDPM gradually add noise to the orginal data $x$ according to a variance schedule $\beta_{1}$, $\beta_{2}$,…,$\beta_{T}$:
$x_{t}=\sqrt{1-\beta_{t}}x_{t-1}+\sqrt{\beta}_{t}\epsilon,\quad t\in\{1,2,...,T-1\}$ (4)
where $\epsilon\sim\mathcal{N}(0,I)$, $x_{0}=x$, $x_{T}\sim\mathcal{N}(0,I)$. During the training, DDPM tends to learn the added noise,
$\mathbb{E}_{\epsilon,x_{0},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},t)\right\|^{2}.$ (5)
The reverse diffusion process gradually denoises standard Gaussian noise into the original data distribution by
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))+\sqrt{\beta_{t}}\epsilon,$ (6)
where $x_{T}\sim\mathcal{N}(0,I)$. After $T$ denoise steps, DDPM can generate the samples from the learned data distribution.
III-B Problem Formulation
The objective of the multi-agent trajectory prediction task is to use the observed history trajectories to jointly generate future trajectories of all the agents in the scene. We denote the accurate and complete history trajectory as $c^{i}=\{S_{-T_{his}+1}^{i},S_{-T_{his}+2}^{i},\dots,S_{0}^{i}\}$ where $S_{k}^{i}$ stands for the two-dimensional position coordinate of $i$th agent at time step $k$. We denote the future trajectory as $y^{i}=\{P_{1}^{i},\dots,P_{T_{fut}}^{i}\},i=1,\dots,N_{a}$. The task is to generate the future trajectory distribution $p(y|c)$. However, the historical trajectory may be missing and injected with noise caused by unexpected conditions such as sensor failure and poor weather. We use an elementwise mask to represent missing trajectory data. We model the injected noise as Gaussian noise so the noisy trajectory is $c+n$, where $n\sim\mathcal{N}(0,\sigma I)$. $\sigma$ is the scale of Gaussian noise.
III-C Represent Full Trajectory Distribution with Diffusion
Previous diffusion-based work [13, 22] formulated it as the conditional prediction process that uses the history trajectory $c$ as the condition to generate the future trajectory distribution $p(y|c)$. Specifically, their denoise module learns to predict the conditional noise value $\epsilon(y_{t},t|c)$. Our key insight here is we can re-formulate the prediction task solving an inverse problem of inpainting, which aims to generate full-length trajectories containing both history and future parts while ensuring the history parts are assembled to the corresponding observed history trajectories. We denote the generated history trajectory as $z_{i}$ and the corresponding full-length trajectory to be $x^{i}=z^{i}\cup y^{i}=\{S_{-T_{his}}^{i},\dots,S_{0}^{i},P_{1}^{i},\dots,P_{T_{fut}}^{i}\}$. We train the DDPM by Eq. (5) where $\epsilon$ has the same shape of $x={x^{1},\dots,x^{N_{a}}}$. Hence we approached to model the unconditional distribution $p(x)$, the one that we can sample from through direct denose steps Eq. (6) at inference time. As the objective remains to be sampling from $p(y|c)$, we first consider adding the conditions by shaping the score-matching process with proper guidance so that we acquire sample from $p(x|c)$. Sequentially, we take the future parts from sampled trajectories by $y=\phi_{fut}(x)$.
III-D Robust Prediction as Posterior Sampling
During inference time, diffusion models sample from the prior distribution by iteratively recovering data from noise, which is a Markovian process described by Eq. (6). For various problems, we have the condition $c$ and the goal is to sample from posterior $p(x|c)$. According to Bayes’ rule: $p(x|c)=p(c|x)p(x)/p(c)$, we have the score of the posterior distribution:
$\nabla_{x}{\log{p_{t}}(x|c)}=\nabla_{x}{\log{p_{t}}(x)}+\nabla_{x}{\log{p_{t}}(c|x)}.$ (7)
According to diffusion posterior sampling (DPS) [16], if the condition or measurement has the form of a general noisy inverse problem $y=\phi(x_{0})+n$, where $\phi$ is an arbitrary operator, and $n\sim\mathcal{N}(0,\sigma^{2}I)$ is the Gaussian noise, then $\nabla_{x}{\log{p_{t}}(c|x)}$ can be approximated as
$\nabla_{x}{\log{p_{t}}(y|x)}=-\lambda\nabla_{x_{t}}\left\|y-\phi(\hat{x_{0}}(x_{t}))\right\|_{2}^{2}=-\lambda\nabla_{x_{t}}\mathcal{L},$ (8)
where $\hat{x_{0}}(x_{t})$ is the posterior mean of $p(x_{0}|x_{t})$:
$\hat{x_{0}}(x_{t})=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(x_{t}+(1-\overline{\alpha}_{t})\epsilon_{\theta}(x_{t},t)).$ (9)
Once the condition operator $\phi$ is known (i.e., $\phi_{his}(x)$ denotes selecting history dimensions from $x$), we can inject such condition by iteratively adding the guidance term Eq. (8) to the intermediate noisy data during the reverse diffusion process so that we drag the sample to the direction that minimizes the guidance loss $\mathcal{L}$. We refer to this characteristic as ”soft-conditioning”, which does not guarantee the exact inpainting of history trajectories. We will show such a design greatly boosts robustness and adaptation abilities in the following sections.
For conventional prediction tasks where we expect data to be clean and complete, soft-conditioning however, may result in ill-conditioned prediction. To further enhance performance on these tasks, we can strictly enforce history conditions following RePaint[24]. This plug-and-play modification introduces an additional step that first compromises observed history trajectories to the noise level $t$ through the forward diffusion process Eq. (4), then concatenates them with the corresponding future parts of intermediate samples.
An implementation of the inference process in our proposed framework is summarized in Algorithm 1.
Algorithm 1 Guided Full Trajectory Diffuser (GFTD)
0:  $\mathcal{L}(\cdot,\cdot)$, $\phi(\cdot)$, $c$, $\{\bar{\alpha}_{t},\beta_{t},\lambda_{t}\}_{t=0}^{T-1}$,
1:  $x_{T}\sim\mathcal{N}(0,\Sigma)$
2:  for $t=T-1,...,1$ do
3:     $\epsilon\sim\mathcal{N}(0,\Sigma)$
4:     $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))+\sqrt{\beta_{t}}\epsilon$
5:     $\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t))$
6:     $g=-\nabla_{x_{t}}\mathcal{L}(\hat{x}_{0},c)$
7:     $x_{t-1}=x_{t-1}+\lambda_{t}g$
8:     if RePaint and $t>0$ then
9:        $\epsilon^{\prime}\sim\mathcal{N}(0,\Sigma)$
10:        $c_{t-1}=\sqrt{1-\beta_{t-1}}c_{0}+\sqrt{\beta}_{t-1}\epsilon$
11:        $\phi_{his}(x_{t-1})=c_{t-1}$
12:     end if
13:  end for
14:  $y=\phi_{fut}(x_{0})$
14:  $y$
III-E Trajectory Latent Representation
Inspired by the advances in image synthesis bought by latent diffusion models [25] and the successful application of latent representation in trajectory prediction [26], we further consider transforming the spatial coordinates to latent space through Principal Component Analysis (PCA). PCA enables a low-rank approximation of trajectory data $X^{i}\in\mathbb{R}^{T*2}$ by mapping it to the latent representation $x^{i}\in\mathbb{R}^{k}$, which stands for linear combination of first-$k$ principal components. This transformation serves to mitigate the effects of noisy data, leading to a more refined and smoother trajectory. Meanwhile, the latent representation of trajectory contains its geometry and temporal characteristics, saving the need for additional temporal encoding steps in the network.
IV Framework Architecture
As we re-formulate the task of prediction as trajectory inpainting, our proposed framework requires a pre-trained DDPM that models the joint distribution of full-length trajectories $p(x)$, which is not conditioned on history motion $c$. Through iteratively adding posterior guidance $-\lambda\nabla_{x_{t}}\mathcal{L}$ during inference time and extract the future part $y$ from generated $x$, we achieve the goal to sample from the conditional distribution $p(y|c)$. We argue that our framework is model-agnostic and that various reasonable network designs of the denoising module are feasible for our methodology. In this section, we present a specific implementation of a lightweight denoising module. We will discuss our data representation method, module architecture, and guidance design in the following.
Data Representation.   To achieve a rotation-invariant representation of agent motion, for each agent in the scene, we rotated and normalized their trajectory to their current position, decoupling the temporal dependency of history motion and social interaction. Parallelly, we extract the spatial relativity of agent trajectories by establishing a graph representation, where node $x^{i},i=1,\dots,N_{a}$ stands for agent motion and edge $E^{ij},i,j=1,\dots,N_{a}$ is characterized by a 6D vector describing their relative position at the current timestep: relative distance $d^{ij}$, direction vector $r^{ij}$ under the reference frame of agent $i$, and relative heading vector $h^{ij}=\{\theta^{ij},cos\theta^{ij},sin\theta^{ij}\}$.
Denoise Module.   Following the standard setting in DDPM, we built a denoise module $\epsilon_{\theta}(x_{t},t,E)$ that predicts the intermediate noise level and denoise the intermediate sample $x_{t}$ by Eq. (6). For each scenario sample, $x_{t}=[PCA(X_{t}^{1}),\dots,PCA(X_{t}^{N_{a}})]$ is the latent node features of agent trajectories at diffusion step $t$. And $E$ stands for the edge information mentioned above, it serves as the only condition of the diffusion model and is shared across all diffusion steps.
The core to $\epsilon_{\theta}$ is a Graph Neural Network (GNN). As depicted in Fig. 2, each latent node feature $x_{t}^{i}$ is firstly concatenated with diffusion time embeddings and then mapped to higher dimensions by a multi-layer perception (MLP). Since node features do not contain global position information, we need to further fuse the nodes with edge features so that we can model the complex interaction. After MLP encoding, the edge is augmented by concatenating with the encoded node
$E_{t}^{ij}=concat\{E^{ij},x_{t}^{i}\},$ (10)
and they are both sent into the Processor which consists of stacked Graph Attention (GAT) [27] layers. Within each GAT layer, nodes are augmented by shared linear transformation parameterized by weight $\bm{W}$. Then, we calculate the attention score by
$\alpha_{t}^{ij}=softmax_{j}(e_{t}^{ij})=\frac{exp(e_{t}^{ij})}{\sum_{k\in N_{a}}exp(e_{t}^{ik})},$ (11)
and
$e_{t}^{ij}=LekyReLU(a(\bm{W}n^{i}||\bm{W}n^{j})),$ (12)
where $a(\cdot)$ linear transformation that maps a vector to a real number. After acquiring interaction information, we map the nodes to the same dimensions as $x_{t}$, resulting in the predicted intermediate noise $\epsilon_{t}$.
Guided Generation.   For trajectory prediction under the description of the noisy inverse problem, the closed-form dependency between the measurement and sample can be formulated as
$c=\phi(X)+n.$ (13)
For the prediction task, $c$ is the observed history trajectories, $\phi$ is the operator that extracts history parts of trajectories from full-length trajectories $X$, and $n$ is the Gaussian noise with adjustable variance based on how clear our history data is. For the original prediction task, we consider the history trajectories to be reliable and the noise level to be low. Similar to the derivation in section $C$, we set our guidance loss as
$\mathcal{L}_{rec}=\left\|c-\phi(\hat{X}_{0}(x_{t}))\right\|_{2},$ (14)
which is the reconstruction loss that measures how precisely our generated trajectories fit the conditions. Inspired by [22], we can ensure a more realistic generation by adding repeller guidance that prevents the agents from colliding with each other:
$\mathcal{L}_{rep}=weight\cdot\frac{1}{N_{a}}\sum_{i,j,k}\mathrm{max}\{(1-\frac{1}{r}dist^{ij}_{k}),0\},$ (15)
where $r$ is the repeller threshold and $dist^{ij}_{k}\in\mathbb{R}^{N_{a}\times N_{a}\times(T_{hist}+T_{fut})}$ is the distance between agent $i$ and $j$ at timestep $t$. Practically, the repeller loss serves as valuable prior knowledge in scenario generation.
V Experiments
In this section, we exhibit the capacity and flexibility of our framework by adapting our model to four distinct tasks without retraining: Trajectory prediction, noisy data prediction, missing data prediction, and controllable generation. We present two models within our proposed framework: the foundational GFTD and its variant, GFTD-RePaint, which is specialized for the basic trajectory prediction task.
Figure 3: Qualitative visualization of controllable generation. Red lines represent the ground truth trajectories. With the given goal point and history, our model can generate goal-oriented trajectories with considerable realism.
TABLE I: Trajectory Prediction Performance on ETH/UCY Dataset. The three best scores are marked by red, blue, and green, respectively. $\dagger$ denotes Joint AgentFormer without the diverse sampler (DLow)
Method minJADE / JFDE (m), K=20 $\downarrow$
ETH(1.4) HOTEL(2.7) UNIV(25.7) ZARA1(3.3) ZARA2(5.9) ETH/UCY Avg.
S-GAN [28] 0.919 / 1.742 0.480 / 0.950 0.744 / 1.573 0.438 / 1.001 0.362 / 0.794 0.589 / 1.212
PECNet [29] 0.618 / 1.097 0.291 / 0.587 0.666 / 1.417 0.408 / 0.896 0.372 / 0.840 0.471 / 0.967
MemoNet[30] 0.499 / 0.859 0.222 / 0.416 0.686 / 1.466 0.349 / 0.723 0.385 / 0.864 0.428 / 0.866
Joint View Vertically [6] 0.652 / 0.839 0.186 / 0.309 0.523 / 1.091 0.331 / 0.634 0.267 / 0.547 0.392 / 0.684
Joint AgentFormer [6] 0.543 / 0.883 0.211 / 0.377 0.596 / 1.247 0.309 / 0.612 0.282 / 0.584 0.388 / 0.741
Ours (GFTD) 0.505 / 0.873 0.174 / 0.297 0.649 / 1.305 0.340 / 0.667 0.308 / 0.620 0.395 / 0.752
Ours (GFTD + RePaint) 0.514 / 0.906 0.191 / 0.329 0.607 / 1.248 0.327 / 0.646 0.288 / 0.585 0.385 / 0.743
TABLE II: Controllable Generation Performance
Method minJADE / JFDE (m), K=20 $\downarrow$
ETH(1.4) HOTEL(2.7) UNIV(25.7) ZARA1(3.3) ZARA2(5.9) ETH/UCY Avg.
Baseline 0.505 / 0.873 0.174 / 0.297 0.649 / 1.305 0.340 / 0.667 0.308 / 0.620 0.395 / 0.752
Goal Point Guidance 0.224 / 0.064 0.069 / 0.033 0.280 / 0.394 0.103 / 0.035 0.094 / 0.032 0.154 / 0.112
V-A Experimental Setups
Datasets.   We carried out all of our experiments on the ETH/UCY [31] [32] dataset, a popular public pedestrian trajectories forecasting benchmark. The dataset contains pedestrian trajectories in bird’s eye view (BEV) from five distinct scenarios (ETH, Hotel, Univ, Zara1, Zara2). We follow the leave-one-out training/evaluation setup that was used in the original S-GAN [28]. Specifically, we partitioned each dataset scene into sliding windows of length 20 steps (8 seconds) at stride 1, with 8 history observation steps (3.2 seconds) followed by 12 prediction steps (4.8 seconds). We then proceed to utilize all sequences containing at least one pedestrian.
Metrics and Baselines.   For scene-level joint prediction, rather than using marginal metrics Average Displacement Error (ADE) and Final Displacement Error (FDE), we follow [6] to use joint metrics JADE/JFDE:
$jointADE(Y,Y^{*})=\frac{1}{TN}\mathop{\mathrm{min}}\limits_{k=1}^{K}\sum\limits_{n=1}^{N}\sum\limits_{t=1}^{T}\left\|s_{t,n}-{s}_{t,n}^{*}\right\|,$ (16)
$jointFDE(Y,Y^{*})=\frac{1}{TN}\mathop{\mathrm{min}}\limits_{k=1}^{K}\sum\limits_{n=1}^{N}\left\|s_{T,n}-{s}_{T,n}^{*}\right\|.$ (17)
The joint metrics emphasize the significance of modeling the interactions between agents by introducing a constraint for top-$K$ evaluations: predictions must originate from the same sample. This approach prevents the overestimation of model performance and penalizes models that fail to accurately simulate the realistic behaviors and motions of all agents present in the scene. In the general prediction task, we compared our model to the state-of-the-art models Joint AgentFormer and Joint View Vertically proposed in [6], as well as other famous baselines in pedestrian prediction: S-GAN [28], PECNet[33], and MemoNet [30]. We compute the best-of-20 JADE/JFDE considering the stochasticity of generative prediction models.
Implementation Details.   The implementation of the denoising module used in experiments contains a three-layer GAT with a hidden size of 128. Additionally, each feed-forward network is realized as a two-layer MLP equipped with post-layer normalization and activated using the Mish activation function [34]. In addition, we observed that training our diffusion model on latent space by Eq. (5) can be unstable and somehow inefficient. Thus, following [35], we modified the DDPM training loss with the Min-SNR-$\gamma$ weighting strategy $\lambda(t)=min\{\gamma,SRN(t)\}/SRN(t)$ to avoid having the model putting too much attention on the final steps of the denoising processing where the noise level is low. We set the training batch size to 32 and we use the Adam optimizer with an initial learning rate of 0.001. All the training is conducted on one GTX-2080Ti GPU.
V-B Trajectory Prediction and Controllable Generation
In the context of basic pedestrian trajectory prediction tasks, we present two models within our Guided Trajectory Diffusion (GFTD) framework: the GFTD model and its augmented version, GFTD with RePaint. To assess their efficacy, we conduct a quantitative evaluation, benchmarking these models against the aforementioned baselines with the ETH/UCY dataset. All models are fairly evaluated by joint metrics to ascertain their optimal scene-level performance across 20 samplings. We directly used the reported baseline performance under joint metrics from [6]. We present the results in TABLE I. In specific for Jonit AgentFormer, we use the pre-trained checkpoint for joint prediction provided by their official GitHub repository (https://github.com/ericaweng/joint-metrics-matter). For a fair comparison, we did not apply the Dlow model for performance enhancement. Our model exhibited relative competitive performance, especially on the HOTEL dataset, where our model outperformed all the baselines.
Controllable trajectory generation is another plausible application of our framework. In addition to the fundamental reconstruction loss Eq. (14), we can incorporate any desirable objectives into the guidance term to control the sampling process. For demonstration, we experimented with goal point generation, i.e., given the ground truth history trajectories and desired goal points $g$, we asked our model to generate future trajectories that reach the goal points. The only modification to adapt our pre-trained model to this particular task is to add an attraction term that measures the L2-norm between the endpoints of generated trajectories and goal points.
We evaluated generation quality through JADE, which represents the realism of the generated scene, and JFDE, which reflects goal-reaching accuracy. As shown in Fig 3, all samples tend to accurately recover the demanded goal point while the generated trajectories are smooth and realistic. The variance of prediction samples tends to be large under cases where the pedestrian performs a sudden turn. However, in most cases, the model successfully covers the modality that most resembles the ground truth. We also present quantitive results in TABLE LABEL:table:_Control, which showcases the effectiveness of goal point guidance.
TABLE III: Prediction with Noisy Data
Noise Std Model minJADE / JFDE (m), K=20 $\downarrow$
ETH(1.4) HOTEL(2.7) UNIV(25.7) ZARA1(3.3) ZARA2(5.9) ETH/UCY Avg.
0.00 Joint AgentFormer 0.543 / 0.883 0.211 / 0.377 0.596 / 1.247 0.309 / 0.612 0.282 / 0.584 0.388 / 0.741
Ours 0.505 / 0.873 0.174 / 0.297 0.649 / 1.305 0.340 / 0.667 0.308 / 0.620 0.395 / 0.752
0.05 Joint AgentFormer 0.588 / 0.992 0.253 / 0.433 0.628 / 1.294 0.362 / 0.708 0.361 / 0.714 0.438 / 0.834
Ours 0.567 / 1.009 0.198 / 0.334 0.703 / 1.388 0.388 / 0.735 0.341 / 0.675 0.439 / 0.828
0.15 Joint AgentFormer 0.765 / 1.287 0.458 / 0.762 0.834 / 1.642 0.649 / 1.254 0.736 / 1.312 0.688 / 1.254
Ours 0.636 / 1.076 0.283 / 0.463 0.872 / 1.661 0.633 / 1.137 0.469 / 0.869 0.579 / 1.041
TABLE IV: Prediction with Missing Data
Missing ratio Model minJADE / JFDE (m), K=20 $\downarrow$
ETH(1.4) HOTEL(2.7) UNIV(25.7) ZARA1(3.3) ZARA2(5.9) ETH/UCY Avg.
25% Joint AgentFormer 2.000 / 3.352 0.610 / 1.043 1.674 / 2.805 1.832 / 3.279 2.093 / 3.092 1.642 / 2.714
Ours 0.528 / 0.912 0.181 / 0.311 0.662 / 1.326 0.342 / 0.677 0.314 / 0.632 0.405 / 0.772
50% Joint AgentFormer 1.763 / 3.007 0.661 / 1.090 1.746 / 2.953 1.887 / 3.379 1.989 / 2.928 1.609 / 2.671
Ours 0.514 / 0.843 0.186 / 0.316 0.703 / 1.390 0.343 / 0.673 0.323 / 0.643 0.414 / 0.773
75% Joint AgentFormer 1.973 / 3.200 0.688 / 1.118 1.936 / 3.347 1.966 / 3.500 1.731 / 2.567 1.659 / 2.746
Ours 0.493 / 0.775 0.241 / 0.423 0.859 / 1.648 0.377 / 0.726 0.419 /0.812 0.478 / 0.877
V-C Prediction with Noisy History
To assess the robustness of our proposed method against noisy input, we perturbed the observed history trajectories by introducing random Gaussian noise $n\sim\mathcal{N}(0,\sigma I)$ of varying standard deviation $\sigma$. Specifically, we introduced two levels of perturbation: a slight level, characterized by a standard Gaussian noise with a 0.05-meter $\sigma$, and a heavy level, with a standard Gaussian noise with a 0.15-meter $\sigma$. Subsequently, we compared the performance of our model with the Joint AgentFormer baseline under identical experimental conditions.
As shown in TABLE LABEL:table:_noisy_data, we observed that our model performs better at heavy noise levels, providing trustworthy prediction even if the data is highly corrupted. The average JADE/JFDE performance of our method on 5 benchmarks is close to the baseline but only degenerated by $46.58\%/38.38\%$ at the heavy level. By contrast, the performance of the baseline method had degenerated by $57.08\%/50.36\%$ at heavy noise level. To reasonably state that the guided posterior sampling accounted for the robustness against noisy data, we further compared it with GFTD-RePaint. GFTD-RePaint includes gradually replacing the history part of the intermediate noisy sample by ground truth history observation, thus ensuring the generated samples were strictly conditioned on the noisy history data. The experimental results in Fig. 4 show that despite having a near performance to GFTD model under clean data conditions, the GFTD-RePaint performs much worse on noisy data, especially under heavy-level noise. This is in accordance with our anticipation, proving the superiority of our proposed framework at noisy conditions rooted in the design of ”soft conditioning” enabled by posterior sampling.
Figure 4: JADE/JFDE performance comparison between GFDT and GFDT+RePaint with noisy data input.
V-D Prediction with Incomplete History
Through the incomplete data prediction task we show the adaptation capacity of our GFTD framework. By masking the initial $25\%$, $50\%$, and $75\%$ steps of history data, we mimic a practical situation where sensors are partially blocked by obstacles and only limited observation is available. We directly evaluated GFTD on these compromised data without retraining. Inferencing missing data could be done easily with no modification to the sampling process since the guidance only enforces conditions using available history steps. As displayed in TABLE LABEL:table:_missing_data, GFTD exhibits competitive performance under incomplete data settings. Even when $75\%$ of history observation is missing, the average JADE/JFDE performance only degenerated for $21.01\%/16.62\%$ and outperformed S-GAN with general data input by a large margin. In comparison, Joint AgentFormer completely failed to present plausible predictions when directly applied to missing data.
VI Conclusion & Discussion
In this work, we present the Guided Full Trajectory Diffuser, a novel diffusion model-based framework for representing the joint distribution of full trajectories, converting trajectory prediction and controllable generation into a unified inverse problem. We formulate the prediction task as spatial-temporal inpainting, a general noisy inverse problem that can be solved through diffusion posterior sampling. Under this framework, the generated trajectories are not rigidly constrained by their historical observations; instead, we approximate such conditions by gradually adding posterior guidance. Such a unique design enables flexibility and adaptivity, resulting in not only competitive performance in joint trajectory prediction tasks but also in controllable generation, particularly in scenarios with noise injection or incomplete historical data. Moreover, our framework is compatible with plug-and-play modules including RePaint, thus expandable to task-oriented enhancements.